ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 141, 2004

Alegre, P., D. E. Blair and A. Carriazo Generalized Sasakianspace-forms

157

Aron, R. and M. Lindström Spectra of weighted composition operators on weighted Banach spaces of analytic functions

263

Bastero, J. and M. Romance A characterization of the MM^{∗}- position of a convex body in terms of covariance matrices

145

Beidar, K. I., M. Brešar, M. A. Chebotar and W. S. Martindale 3-rd Polynomial preserving maps on certain Jordan algebras

285

Bialy, M. L. and R. S. MacKay Symplectic twist maps without conjugate points

235

Blair, D. E. See P. Alegre

Borichev, A. Slow area-preserving diffeomorphisms of the torus  277

Brešar, M. See K. I. Beidar

Carriazo, A. See P. Alegre

Chebotar, M. A. See K. I. Beidar

Galindo, J. and M. Sanchis Stone–Weierstrass theorems for groupvalued functions

341

Gillespie, T. A. and J. L. Torrea Dimension free estimates for the oscillation of Riesz transforms

125

Ivanov, S. V. and M. G. Katz Generalized degree and optimal Loewner-type inequalities

221

Katz, M. G. See S. V. Ivanov

Lindström, M. See R. Aron

Lucchini, A. A 2-generated just-infinite profinite group which is not positively generated

119

MacKay, R. S. See M. L. Bialy

Magaard, K. and H. Völklein The monodromy group of a function on a general curve

355

Martindale 3-rd, W. S. See K. I. Beidar

Mildenberger, H. and S. Shelah On needed reals  1

Müller, D. and A. Seeger Singular spherical maximal operators on a class of two step nilpotent Lie groups 315

Pablos Romo, F. A generalization of the Contou-Carrère symbol  39

Romance, M. See J. Bastero Roquette, P. On the local-global principle for embedding problems

over global fields

369

Rørdam, M. A purely infinite AH-algebra and an application to AF-embeddability

61

Rubin, B. Reconstruction of functions from their integrals over

k-planes 93

Sanchis, M. See J. Galindo Seeger, A. See D. Müller Shelah, S. See H. Mildenberger Srebro, U. and E. Yakubov Boundary behavior of μ-homeo-

morphisms

249

Tent, K. Half-Moufang generalized hexagons  83

Torrea, J. L. See T. A. Gillespie

Vestfrid, I. A. Affine properties and injectivity of quasi-isometries  185

Völklein, H. See K. Magaard Yakubov, E. See U. Srebro

Yu, H. On Tate–Shafarevich groups over Galois extensions  211